Assigns

Version 1.1 February,14 1993

by Enrico Fedrigo

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

We don’t guarantee that the programs described in this document are 100% reliable. You are using this material at your own risk. The author can not be made responsible for any damage which is caused by using these programs.

This package is freely distributable, but still copyright by Enrico Fedrigo. This means that you can copy it freely as long as you don’t ask for a more than nominal copying fee. This fee must not be more than US $5.

Permission is granted to include this package in Public-Domain collections, especially in Fred Fish Amiga Disk Library (including CD ROM versions of it). The distribution file may be uploaded to Bulletin Board Systems or FTP servers. If you want to distribute this program you must use the original distribution archive ‘Assigns1_0.lha’.

1 Address

Author may be reached at:

Postal Address

                   Enrico Fedrigo
                   via Giacomo Leopardi, 6
                   31033 Castelfranco Veneto (Treviso)
                         ITALY

Electronic Mail (InterNet):

               fedro@sabrina.dei.unipd.it
               fedro@paola.dei.unipd.it

2 Introduction

Assigns is a preference editor that fills a lack of WorkBench.

Using Assigns you don’t need any more tens of assign lines in your “User-Startup”, it’s all made in a single line (or even less, see Chapter 3 [Installation], page 4).

Assigns is not an “Assign-Manager”, it has to be used as an utility to start when booting the system, and it handles all your assigns. It can even add assigns and activate them just as c:assign (but these cannot be removed).

I’m thinking about adding assign removal, to make of Assigns a complete independent substitutive utility, but this operation seems to be discouraged even by the operating system manual (see Chapter 9 [Future develops], page 11).

About this we may note that doesn’t exists a system call meant to remove assignments, neither DEFER or PATH, and then there isn’t a “legal” way to remove them.

Assigns is the first program I made that I decided to leave of public domain, and I’ve been careful following all Commodore standards. Thus the window is an AppWindow font-sensitive, it works with iff files, accepts the command “version”, can be used either from WorkBench or from CLI, and uses the same kind of menus of all the other prefs.

3 Installation

The installation is really easy:

− Copy “Assigns” and “Assigns.info” in “SYS:prefs”. If your processor is a 68030 or 68040 you may choose the optimized version named “Assigns.030”. Copy and rename it in place of normal “Assigns”.

− Copy in LOCALE:Catalogs/<language> the file of your language, if is present in the archive. If not, you can fill the empty catalog Assigns.ct: if you will send me, I’ll include it in the next version.

− Activate Assigns − use option “Read System List” from “Edit” menu − Delete with “Del” button all unmodified default system assignments in the assign list, such as:

        C: CLIPS: DEVS: ENV: ENVARC: FONTS:
        HELP: KEYMAPS: L: LIBS: LOCALE:
        PRINTERS: REXX: S: SYS: T:

− Did you add custom assigns in your User-Startup or Startup-Sequence?

• If yes: − Save your list in “SYS:WbStartup” with “Save As...” from “Project” menu

• If no: − Save configuration with “SAVE” button − Add before the first call of your custom assign:

               SYS:Prefs/Assigns USE

The installation is completed.

4 Use

Assign has either CLI and WB interface. Activating the “EDIT” flag or executing Assigns from WorkBench editing window is opened. Opening the program by icon you can set some tooltypes similar to CLI parameters. Of course

these tooltypes may be selected even in the project icons created by Assigns and used to load it.

To add a new assignment just click on ‘ADD’ gadget of the Assign-List. You’ll be asked for a new Target.

If desired target directory doesn’t exists yet, exit from the file requester with “CANCEL”. This will activate the gadget string for input.

You may choose one of the following types:

             ABSOLUTE
             DEFER
             PATH

You may want to know whether or not an assignment is succesfully activated or not (WARN) Last you may need to activate immediately the assign, selecting “IMMEDIATE” button. This

may be useful to add assigns referred to other assignments: Assigns handles fine even these. If you choose “ABSOLUTE” type you will can add more TARGETS by clicking on “ADD” button. “DEL” button delete the selected target.

Assigns allows even to add targets by choosing the assign from the assign-list and then carry the path (directory o device) icon above Assigns window.

This is useful to select several targets simultaneously. If the icon was a valid preference file, it’ll be used. By a single click on a TARGET you can modify its path using the gadget string. With a double

click you choose it from a file requester.

From menu you can choose other options.

The last three buttons: •‘SAVE’ Saves configuration on disk, executes the assignments and exits. •‘USE’ Saves configuration on RAM:, executes the assignments and exits. The configuration

won’t be available after a reboot.

•‘CANCEL’ Exits without saving or executing assignments.

Obviously, as any preference editor, even Assigns may be started once a time: starting a second instance returns only a “popup” of the existent Assigns window and its activation.

5 CLI Execution

CLI syntax (available typing ‘Assigns ?’):

FROM,EDIT/S,USE/S,SAVE/S,PUBSCREEN/K,BOX3D/S

−‘FROM’ File to be load instead of “ENV:Sys/Assign.prefs”. FROM key may be

omitted.

−‘EDIT’ [Default]

Once loaded prefs file, opens editing window to allow changes

−‘USE’ Once loaded prefs file, executes assignments and exits. −‘SAVE’ Once loaded prefs file, saves it on “ENVARC:Sys/Assign.prefs”, executes

the assignments and exits.

−‘PUBSCREEN’ <pub screen name>

Opens editing window on indicated public screen.

−‘BOX3D’ Around the two groups of gadgets “Assign” and “Target(s)” there is a NeXT-like box. Selecting Yes or No you choose to use or not it.

6 ToolTypes

The availables ToolTypes (usable even by Assigns’s icons) are:

−‘EDIT’ [Default]

Once loaded prefs file, opens editing window to allow changes

−‘USE’ Once loaded prefs file, executes assignments and exits. −‘SAVE’ Once loaded prefs file, saves it on “ENVARC:Sys/Assign.prefs”, executes

the assignments and exits.

−‘CREATEICONS=YES|NO’ [Default = YES] Choose Yes/No whether or not you want to create icons for the files created with ’Save As...’ from menu (see Chapter 7 [Menu], page 9).

−‘PUBSCREEN’ <pub screen name>

Opens editing window on indicated public screen.

−‘BOX3D=YES|NO’ [Default = YES] Around the two groups of gadgets “Assign” and “Target(s)” there is a NeXT-like box. Selecting Yes or No you choose to use or not it.

You may select EDIT,USE and SAVE even as:

ACTION=EDIT | USE | SAVE

7 Menu

•‘Project’ −‘Open...’ Allows to load a previously saved prefs file. −‘Save As...’

Saves assign-lists. If set CREATEICONS YES, an icon is added to the file. The icon may be used to activate the assign-list without starting Assigns. Default tooltype is ‘AC-TION=USE’.

−‘Quit’ Quits the program.

•‘Edit’

−‘Read System List’ Reads all present system or custom assignments. You lose the currents eventually listed assignments.

−‘Last Saved’ Reads last saved assignments from ENVARC: You lose the currents eventually listed assignments.

−‘Restore’ Loads last valid assignments from ENV: You lose the currents eventually listed assignments.

•‘Settings’

−‘Create Icons?’ (Flag) Tells whether or not to add an icon to the file generated by Assigns with “Save As...” option. When starting takes tooltype (see Chapter 6 [ToolTypes], page 8) CREATEICONS value. Default is Yes.

8 Distribution

The complete distributed Assigns package includes:

Assigns1.1.info
Assigns1.1/Assigns  • 680X0 version
Assigns1.1/Assigns.030  • 68030/881 version
Assigns1.1/Assigns.info
Assigns1.1/Docs.info
Assigns1.1/Docs/Italiano.doc
Assigns1.1/Docs/Italiano.doc.info
Assigns1.1/Docs/Italiano.guide
Assigns1.1/Docs/Italiano.guide.info
Assigns1.1/Docs/Italiano.dvi
Assigns1.1/Docs/Italiano.dvi.info
Assigns1.1/Docs/Italiano.texi
Assigns1.1/Docs/Italiano.texi.info
Assigns1.1/Docs/English.doc  • plain ASCII docs
Assigns1.1/Docs/English.doc.info
Assigns1.1/Docs/English.guide  • AmigaGuide docs
Assigns1.1/Docs/English.guide.info
Assigns1.1/Docs/English.dvi  • DVI docs
Assigns1.1/Docs/English.dvi.info
Assigns1.1/Docs/English.texi  • TeXinfo sources
Assigns1.1/Docs/English.texi.info
Assigns1.1/Locale/Catalogs/Italiano/Assigns.catalog
Assigns1.1/Locale/Catalogs/Assigns.ct

These files MUST be present.

9 Future develops

Now the program works fine and is reliable, next version will be optimized in memory usage. I intend to write the next version of Assigns in C++ and I’m beginning to study some objects

for programming in AmigaDOS.

Lastly: make of Assigns an “Assign-Manager”, but there are some problems.

Chapter 10: Versions 12

10 Versions

Assigns1.0: first release, full working. Only for betatesters

Assigns1.1: first public release. Changes: localization and some bugs fixed, especially the wrong activation of some gadgets.

11 Greetings

My greetings for the help (tests and suggestions) are for:^{1}:

Alessandro Borille    paride@sabrina.dei.unipd.it
Eugenio Castellani    aj2@sabrina.dei.unipd.it
Fabio D'Alessi        phobia@sabrina.dei.unipd.it
Sergio Fantinel       ik3fan@paola.dei.unipd.it
Jacopo Farinato       redi@sabrina.dei.unipd.it
Gianluigi Nusca       gigio@paola.dei.unipd.it
Emanuele Verza        alien@sabrina.dei.unipd.it

Specially Alessandro and Eugenio for italian to english translation.

1 alphabetically listed